OpenStudy (anonymous):

When comparing 22/42 to 1/2 using mental math, what fraction is useful?

OpenStudy (anonymous):

21/42 ?

OpenStudy (anonymous):

\[\frac{21}{42}=\frac{1}{2}*\frac{21}{21}=\frac{1}{2}\]

OpenStudy (anonymous):

Thank you!